OpenStudy (anonymous):

1-3 cos= sin? need help getting started with my hw :/

OpenStudy (rogue):

Are you sure that is the whole problem? Your missing some variables :P

OpenStudy (anonymous):

yea the only thing im missing is the thetas which arent on the keyboard lol

OpenStudy (rogue):

Any cosine or sine squareds? Its not possible to solve that one by hand unless there's a power in it.

OpenStudy (anonymous):

nope u have to replace sin with 1-cossquared useing identities idk where to go from there

OpenStudy (rogue):

Oh, okay. Is this your problem?\[1-3 \cos \theta = \sin^2 \theta\]